OpenStudy (anonymous):

What is ln e^-6? please with steps!

OpenStudy (anonymous):

1/e^6. an that can be approximated.

geerky42 (geerky42):

There is property of logarithm where \(\log(x^n) = n\log(x)\) Try apply that property.

geerky42 (geerky42):

You should have \(-6\cdot\ln(e)\)

OpenStudy (anonymous):

where do you see a log here though? And it is not stated to be equal to anything (not equal to any other side of the equation... and there is no equation)... you know what I mean?

geerky42 (geerky42):

Looks like you need help with this question too.

geerky42 (geerky42):

From this property , we have \(\ln(e^{-6}) ~~~\Longleftrightarrow~~~-6\ln(e)\) And I'm sure you know what \(\ln(e)\) is, do you? @coconuttree

OpenStudy (anonymous):

yes 1

geerky42 (geerky42):

Right, so we have \(-6\cdot1 = \boxed{-6}\) Makes sense so far?

OpenStudy (anonymous):

yes very much so

OpenStudy (anonymous):

I didn't see the "ln" of e^(-6). sorry.

OpenStudy (anonymous):

I though ti is just e^(-6).

OpenStudy (anonymous):

soury